A Complete Bibliography of Distributed Ledger Technologies: Research and Practice (DLT)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

01 July 2023 Version 1.00

Title word cross-reference

ACM [CH22]. Activities [ATS22]. Adversarial [ATS22, RYPD23]. Algorand [Dim22]. Analysis [WLCX23]. Analyzing [ATS22]. aware [TVWT22].

Based [CGK23, STZS22, WLW^{+}23, ATS22]. Behavior [ATS22]. Block [RDK23]. Blockchain [ATVK22, BZHS23, KXZ^{+}23, RYPD23, SSTV23, SLS^{+}22, STZS22, WLW^{+}23]. Blockchain-Based [STZS22]. Blockchains [AAA22, ATS22, SDZ23, SMB23]. BlockQoS [SDZ23]. Business [BZHS23]. Byzantine [CGK23, Wan22].

Cardossier [STZS22]. Case [SSTV23, STZS22]. Casper [GLMV23]. chain [HYD^{+}23, AAA22]. Chain-Net [AAA22]. Checking [BADR^{+}23]. Cloud [BADR^{+}23]. Collaborative [WTRM22]. Compliance [BADR^{+}23]. Concurrency [SLS^{+}22]. Consensus [ARP23]. Consortia [BZHS23]. Container [WTRM22]. Contracts [ZSFD23]. Control [TVWT22]. Cross [HYD^{+}23]. Cross-chain [HYD^{+}23]. Crowdsensing [CGK23]. Cryptocurrencies [ARP23, SMB23]. Cryptocurrency [ATS22]. Cyber [RYPD23].

Data [KXZ^{+}23, RDK23, SSTV23, STZS22, ZSFD23]. dBFT [WLCX23]. Decentralization [TVWT22].

Decentralized [QWCX22, ZSFD23]. demand [SDZ23]. Design [BADR^{+}23]. Detecting [ATS22]. Digital [SSTV23, STZS22]. Distributed [BRH^{+}23, CH22, RDK23, Wan22]. DLTs [TVWT22]. Do [BRH^{+}23]. Double [SSTV23]. Double-spending [SSTV23]. Dual [SLS^{+}22]. Dual-Blockchain [SLS^{+}22].

E-prescription [SSTV23]. Editorial [CH22]. Encryption [WLW^{+}23]. Engage [BZHS23]. Ethereum [ATS22, RYPD23]. Exchange [SSTV23]. Extending [RDK23].

Fair [SDZ23]. Fault [CGK23, Wan22]. Federated [QWCX22]. Flow [TVWT22]. Formal [WLCX23]. Framework [AAA22, ARP23].

General [SMB23]. General-purpose [SMB23]. Generative [RYPD23]. Genomics [ATVK22].

Harmonizing [SSTV23]. Healthor [TVWT22]. Heterogeneity [TVWT22]. Heterogeneity-aware [TVWT22]. Hunting [RYPD23]. Hybrid [GLMV23, QWCX22]. Hyperledger [RDK23].

Implementation [BADR^{+}23, RDK23]. Inaugural [CH22]. Incentive [QWCX22]. Incentivized [ZSFD23]. Incentivizing [STZS22]. Increase [TVWT22]. InDaMul [ZSFD23]. Infrastructure [WTRM22]. inspired [AAA22]. Internet [AAA22]. Internet-inspired [AAA22]. Interoperability [BRH^{+}23]. Interoperable [AAA22]. Issue [CH22].

Learning [QWCX22]. Ledger [BRH^{+}23, CH22, RDK23]. Ledgers [Wan22]. Legal [ARP23]. Literature [ATVK22].

Malicious [ATS22]. Management

[SSTV23]. Marketplace [KXZ^{+}23]. Matrix [RDK23]. Mechanism [QWCX22]. Members [BZHS23]. Mobile [CGK23]. Monetization [SDZ23]. Motives [BZHS23]. Mules [ZSFD23].

Need [BRH^{+}23]. NEO [WLCX23]. Net [AAA22]. Networking [ZSFD23]. Networks [RYPD23]. NimbleChain [SMB23].

On-demand [SDZ23]. Opportunistic [ZSFD23]. Optimal [ARP23]. Optimally [SLS^{+}22].

Parameters [GLMV23]. Performance [TVWT22]. Permissionless [ATS22, SMB23]. Perspective [BZHS23]. Platform [KXZ^{+}23]. Practical [CGK23]. Practice [CH22]. prescription [SSTV23]. Preserving [KXZ^{+}23]. Prevention [SSTV23]. Privacy [KXZ^{+}23, RDK23]. Privacy-Preserving [KXZ^{+}23]. Proof [Dim22]. Proof-of-Stake [Dim22]. Protocol [WLCX23]. Providers [BADR^{+}23]. purpose [SMB23].

Quality [SDZ23, STZS22]. Quality-of-Service [SDZ23].

Registration [WLW^{+}23].

Registration-Based [WLW^{+}23]. Regulate [ARP23]. Reinshard [SLS^{+}22]. Repositories [WTRM22]. Requirements [RDK23]. Research [CH22]. Resilience [GLMV23]. Resolution [SLS^{+}22]. Review [ATVK22]. Revisited [Wan22]. Robustness [CGK23].

Scalable [WTRM22]. Security [WLCX23]. Sensitive [SSTV23]. Service [SDZ23].

Sharded [SLS^{+}22]. Should [ARP23]. Smart [ZSFD23]. Socio [BZHS23]. Socio-technical [BZHS23]. Solution [BRH^{+}23]. Speeding [SMB23]. spending [SSTV23]. Stake [Dim22]. Strategic [ARP23]. Survey [HYD^{+}23]. Systematic [ATVK22]. Systems [STZS22, ZSFD23].

technical [BZHS23]. Technologies [CH22, HYD^{+}23]. Technology [BZHS23, BRH^{+}23, RDK23]. Threat [RYPD23]. Throughput [ARP23]. Tolerance [CGK23, Wan22]. Transaction [ARP23]. Transparent [WLW^{+}23]. Trustworthy [WTRM22].

Usecase [ATS22]. using [SDZ23].

Values [GLMV23]. Varying [GLMV23]. via [ARP23].

Wallets [SSTV23].

References

Abdullah:2022:CNI

[AAA22] Sidrah Abdullah, Junaid Arshad, and Muhammad Alsadi. Chain-Net: an Internetinspired framework for interoperable blockchains. Distributed Ledger Technologies: Research and Practice (DLT), 1(2):7:1– 7:??, December 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3554761.

Ahuja:2023:HSW

[ARP23] Aditya Ahuja, Vinay Ribeiro, and Ranjan Pal. How should we regulate cryptocurrencies via consensus?: A strategic framework for optimal legal transaction throughput. Distributed Ledger Technologies: Research and Practice (DLT), 2(1):4:1– 4:??, March 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3567593.

Agarwal:2022:AMA

[ATS22] Rachit Agarwal, Tanmay Thapliyal, and Sandeep Shukla. Analyzing malicious activities and detecting adversarial behavior in cryptocurrency based permissionless blockchains: an Ethereum usecase. Distributed Ledger Technologies: Research and Practice (DLT), 1(2):8:1– 8:??, December 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3549527.

Alghazwi:2022:BGS

[ATVK22] Mohammed Alghazwi, Fatih Turkmen, Joeri Van Der Velde, and Dimka Karastoyanova. Blockchain for genomics: a systematic literature review. Distributed Ledger Technologies: Research and Practice (DLT), 1(2):11:1– 11:??, December 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3563044.

Barati:2023:CCC

[BADR^{+}23] Masoud Barati, Kwabena Adu-Duodu, Omer Rana, Gagangeet Singh Aujla, and Rajiv Ranjan. Compliance checking of cloud providers: Design and implementation. Distributed Ledger Technologies: Research and Practice (DLT), 2(2):13:1–13:??, June 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3585538.

Belchior:2023:DYN

[BRH^{+}23] Rafael Belchior, Luke Riley, Thomas Hardjono, André Vasconcelos, and Miguel Correia. Do you need a distributed ledger technology interoperability solution? Distributed Ledger Technologies: Research and Practice (DLT), 2(1):1:1–1:??, March 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3564532.

Bauer:2023:WBS

[BZHS23] Ingrid Bauer, Rafael Ziolkowski, Janine Hacker, and Gerhard Schwabe. Why blockchain: a socio-technical perspective on the motives of business consortia members to engage with blockchain technology. Distributed Ledger Technologies: Research and Practice (DLT), 2(2): 10:1–10:??, June 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3573893.

Chen:2023:PBF

[CGK23] Zhiyan Chen, Omer Melih Gul, and Burak Kantarci. Practical Byzantine fault tolerance based robustness for mobile crowdsensing. Distributed Ledger Technologies: Research and Practice (DLT), 2(2):12:1–12:??, June 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3580392.

Choo:2022:EII

[CH22] Kim-Kwang Raymond Choo and Mohammad Hammoudeh. Editorial: The inaugural issue of ACM Distributed Ledger Technologies: Research and Practice. Distributed Ledger Technologies: Research and Practice (DLT), 1(1):1:1–1:??, September 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3559010.

Dimitri:2022:PSA

[Dim22]  Nicola Dimitri.  Proof-of-

stake in Algorand. Distributed Ledger Technologies: Research and Practice (DLT), 1(2):9:1– 9:??, December 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3550197.

Galletta:2023:RHC

[GLMV23] Letterio Galletta, Cosimo Laneve, Ivan Mercanti, and Adele Veschetti. Resilience of hybrid Casper under varying values of parameters. Distributed Ledger Technologies: Research and Practice (DLT), 2(1):5:1– 5:??, March 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3571587.

Han:2023:SCC

[HYD^{+}23] Panpan Han, Zheng Yan, Wenxiu Ding, Shufan Fei, and Zhiguo Wan. A survey on crosschain technologies. Distributed Ledger Technologies: Research and Practice (DLT), 2(2):15:1– 15:??, June 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3573896.

Klaine:2023:PPB

[KXZ^{+}23] Paulo Valente Klaine, Hao Xu, Lei Zhang, Muhammad Imran, and Ziming Zhu. A privacypreserving blockchain platform for a data marketplace. Distributed Ledger Technologies: Research and Practice (DLT), 2 (1):7:1–7:??, March 2023. CO-DEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3573894.

Qi:2022:HIM

[QWCX22] Minfeng Qi, Ziyuan Wang, Shiping Chen, and Yang Xiang. A hybrid incentive mechanism for decentralized federated learning. Distributed Ledger Technologies: Research and Practice (DLT), 1(1):4:1–4:??, September 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3538226.

Roberts:2023:DBM

[RDK23] Joshua D. Roberts, Joanna F. Defranco, and D. Richard Kuhn. Data block matrix and hyperledger implementation: Extending distributed ledger technology for privacy requirements. Distributed Ledger Technologies: Research and Practice (DLT), 2 (2):16:1–16:??, June 2023. CO-DEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3585539.

Rabieinejad:2023:GAN

[RYPD23] Elnaz Rabieinejad, Abbas Yazdinejad, Reza M. Parizi, and Ali Dehghantanha. Generative adversarial networks for cyber threat hunting in Ethereum blockchain. Distributed Ledger Technologies: Research and Practice (DLT), 2 (2):9:1–9:??, June 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3584666.

Shabir:2023:BFM

[SDZ23] Muhammad Muneem Shabir, Syed Muhammad Danish, and Kaiwen Zhang. BlockQoS: Fair monetization of on-demand quality-of-service using blockchains. Distributed Ledger Technologies: Research and Practice (DLT), 2 (2):11:1–11:??, June 2023. CO-DEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3580284.

Sharma:2022:ROS

[SLS^{+}22] Vishal Sharma, Zengpeng Li, Paweł Szałachowski, Teik Guan Tan, and Jianying Zhou. Reinshard: an optimally sharded dual-blockchain for concurrency resolution. Distributed Ledger Technologies: Research and Practice (DLT), 1(1):5:1–5:??, September 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3547300.

Silva:2023:NSC

[SMB23] Paulo Silva, Miguel Matos, and João Barreto. NimbleChain: Speeding up cryptocurrencies in general-purpose permissionless blockchains. Distributed Ledger Technologies: Research and Practice (DLT), 2(1):8:1– 8:??, March 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3573895.

Schlatt:2023:HSD

[SSTV23] Vincent Schlatt, Johannes Sedlmeir, Janina Traue, and Fabiane Völter. Harmonizing sensitive data exchange and doublespending prevention through blockchain and digital wallets: The case of E-prescription management. Distributed Ledger Technologies: Research and Practice (DLT), 2(1):6:1–6:??, March 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL https://dl.

                                        acm.org/doi/10.1145/3571509.

Spychiger:2022:IDQ

[STZS22] Florian Spychiger, Claudio J. Tessone, Liudmila Zavolokina, and Gerhard Schwabe. Incentivizing data quality in blockchain-based systems — the case of the digital cardossier. Distributed Ledger Technologies: Research and Practice (DLT), 1(1):3:1–3:??, September 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3538228.

Theis:2022:HHA

[TVWT22] Jonas Theis, Luigi Vigneri, Lin Wang, and Animesh Trivedi. Healthor: Heterogeneity-aware flow control in DLTs to increase performance and decentralization. Distributed Ledger Technologies: Research and Practice (DLT), 1(2):10:1–10:??, December 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3555676.

Wang:2022:BFT

[Wan22]  Yongge Wang.  Byzantine

fault tolerance for distributed ledgers revisited. Distributed Ledger Technologies: Research and Practice (DLT), 1(1):2:1– 2:??, September 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3538227.

Wang:2023:FSA

[WLCX23] Qin Wang, Rujia Li, Shiping Chen, and Yang Xiang. Formal security analysis on dBFT protocol of NEO. Distributed Ledger Technologies: Research and Practice (DLT), 2(1):2:1– 2:??, March 2023. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (electronic). URL

https://dl.acm.org/doi/10.
1145/3568314.

Wang:2023:TRB

[WLW^{+}23] Qin Wang, Rujia Li, Qi Wang, David Galindo, Shiping Chen, and Yang Xiang. Transparent registration-based encryption through blockchain. Distributed Ledger Technologies: Research and Practice (DLT), 2 (1):3:1–3:??, March 2023. CO-DEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3568315.

Wei:2022:STI

[WTRM22] Franklin Wei, Stephen Tate, Mahalingam Ramkumar, and Somya Mohanty. A scalable trustworthy infrastructure for collaborative container repositories. Distributed Ledger Technologies: Research and Practice (DLT), 1(1):6:1–6:??, September 2022. CODEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3554760.

Zichichi:2023:IID

[ZSFD23] Mirko Zichichi, Luca Serena, Stefano Ferretti, and Gabriele D’angelo. InDaMul: Incentivized data mules for opportunistic networking through smart contracts and decentralized systems. Distributed Ledger Technologies: Research and Practice (DLT), 2 (2):14:1–14:??, June 2023. CO-DEN ???? ISSN 2769-6480 (print), 2769-6472 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3587696.